the rate of the jet stream is 300 miles per hour. traveling with the jet stream, an airplane can fly 2500 miles in the same amount of time as it takes to fly 1000 miles against the jet stream. What is the airplanes average rate in calm air?
Answer by rwm(914)

(a+300)*t=2500
(a-300)*t=1000
a=700
t=2.5